Could someone help me?
1.)Solve the system of linear equations by using matrices: [9,3,4] [x1] [7
[ 4,3,4] [x2] = [8
[1,1,1] [x3] [3
2. If logk4=2/3 then what does k = ?


Answer by venugopalramana(3286)

Solve the system of linear equations by using the matrices: [9,3,4][x1]=7
[4,3,4][x2]=8
[1,1,1][x3]=3
9 3 4 7
4 3 4 8
1 1 1 3
R1=R1/9
1 0.333333333 0.444444444 0.777777778
4 3 4 8
1 1 1 3
R2=R2-4R1…..R3=R3-R1
1 0.333333333 0.444444444 0.777777778
0 1.666666667 2.222222222 4.888888889
0 0.666666667 0.555555556 2.222222222
R2=R2/(5/3)
1 0.333333333 0.444444444 0.777777778
0 1 1.333333333 2.933333333
0 0.666666667 0.555555556 2.222222222
R3=R3-R2*2/3
1 0.333333333 0.444444444 0.777777778
0 1 1.333333333 2.933333333
0 0 -0.333333333 0.266666667
R3=-R3*3
1 0.333333333 0.444444444 0.777777778
0 1 1.333333333 2.933333333
0 0 1 -0.8
R2=R2-R3*4/3….R1=R1-R3*4/9
1 0.333333333 0 1.133333333
0 1 0 4
0 0 1 -0.8
R1=R1-R2/3
1 0 0 -0.2
0 1 0 4
0 0 1 -0.8
HENCE X1=-0.2 X2=4 X3=-0.8